### v2

@@ Abstract
This paper presents the first application of Gaussian Mixture Copula Models to the statistical modeling of driving scenarios for the safety validation of automated driving systems. Knowledge of the joint probability distribution of scenario parameters is essential for scenario-based safety assessment, where risk quantification depends on the likelihood of concrete parameter combinations. Gaussian Mixture Copula Models bring together the multimodal expressivity of Gaussian Mixture Models and the flexibility of copulas, enabling separate modeling of marginal distributions and dependencies. We benchmark Gaussian Mixture Copula Models against previously proposed approaches—Gaussian Mixture Models and Gaussian Copula Models—using real-world driving data drawn from scenarios defined in United Nations Regulation No. 157. Our evaluation across approximately 18 million scenario instances demonstrates that Gaussian Mixture Copula Models consistently surpass Gaussian Copula Models and perform better than, or at least comparably to, Gaussian Mixture Models, as measured by both log-likelihood and Sinkhorn distance. These results are promising for the adoption of Gaussian Mixture Copula Models as a statistical foundation for future scenario-based validation frameworks.
@@ 1 INTRODUCTION
  • Gaussian Mixture Models (GMMs) approximate the joint density with a weighted sum of elliptical Gaussians [14] and have been successfully applied to driving scenarios, such as the highway cut-in risk analysis by Akagi et al. [15]. Nevertheless, for finite mixtures, GMMs can be limited in practice at capturing intricate dependencies between parameters due to their reliance on Gaussian assumptions. These assumptions often oversimplify complex relationships and higher-order interactions, restricting their ability to model non-linear dependencies effectively [16].
@@ A GMM describes the joint density of random variables as a weighted sum of $K$ normal distributions,
where $\alpha^{k}\geq 0$ denotes the weight of the $k$-th component, such that $\sum_{k=1}^{K}\alpha^{k}=1$. $\phi(x_{1},x_{2},\dots x_{d};\bm{\mu}^{k},\Sigma^{k})$ represents a multivariate normal distribution with mean vector $\bm{\mu}^{k}$ and covariance matrix $\Sigma^{k}$. The parameters $\Theta=\{\alpha^{k},\bm{\mu}^{k},\Sigma^{k}\}_{k=1}^{K}$ of a GMM can be efficiently learned by using the expectation-maximization (EM) algorithm [24]. The two-step approach of the EM algorithm guarantees convergence to a local maximum of the likelihood function. As stated before, a major limitation of GMMs is the assumption that each of the $K$ components follows a multivariate normal distribution. This assumption can become restrictive in practice when the data exhibit non-Gaussian characteristics, such as skewness, heavy-tailed distributions, or intricate dependencies between variables, like this is the case for our scenario parameters.
@@ 3 ESTIMATION OF GMCM PARAMETERS
Estimating the dependencies among the random variables, represented by the GMC parameters $\Theta=\left\{\alpha^{k},\mkern 1.0mu\bm{\mu}^{k},\mkern 1.0mu\Sigma^{k}\right\}_{k=1}^{K}$ is non-trivial. Neither it is possible to directly determine these from a empirical correlation matrix, like this is the case for Gaussian copulas. Nor the EM algorithm for GMM parameter estimation is directly applicable. The (log-)likelihood function (7) of GMCMs does not admit a closed analytical form, in particular because of function $\Psi^{-1}$ appearing in it. Further, GMCMs suffer from an inherent issue of parameter non-identifiability. These challenges and structural properties have been previously examined by Tewari et al. [19, 23].
@@ 4 EXPERIMENT
In this section, we present our experimental study of GMCMs for joint probability estimation of scenario parameters. We demonstrate that GMCMs surpass GCMs and perform better than, or at least comparably to, GMMs in accurately capturing the complex parameter dependencies. Our experiments conducted on real‐world datasets for certification relevant scenarios, confirm the previously reported advantages of GMCMs [19, 23]. They also extend these benefits to the domain of scenario parameter modeling.
@@ 4.1 REAL-WORLD SCENARIO DATASETS
The scenario Deceleration is provided as one of three traffic critical disturbance scenarios to define conditions under which ALKS shall avoid a collision. The conditions, respectively parameters of the scenarios, are determined by simulating these scenarios with an attentive human driver performance model.
@@ 4.2 RESULTS
To illustrate the achieved expressive power of the GMCM, we visualize the bivariate marginal distribution of LeadVehicle Init HeadwayTime (s) and Ego InitSpeed Ve0 (kph) from the Follow Lead Vehicle Emergency Brake dataset via density heatmaps. Fig. 4 compares the original data distribution—which appears multimodal with non-Gaussian modes—with the estimated densities from the GMCM, the GMM, and the GCM. Visual inspection of the bivariate density heatmaps suggests that the GMCM most accurately captures the real data distribution.

### v3

@@ Abstract
This paper presents the first application of Gaussian Mixture Copula Models to the statistical modeling of driving scenarios for the safety validation of automated driving systems. Knowledge of the joint probability distribution of scenario parameters is essential for scenario-based safety assessment, where risk quantification depends on the likelihood of concrete parameter combinations. Gaussian Mixture Copula Models bring together the multimodal expressivity of Gaussian Mixture Models and the flexibility of copulas, enabling separate modeling of marginal distributions and dependencies. We benchmark Gaussian Mixture Copula Models against previously proposed approaches—Gaussian Mixture Models and Gaussian Copula Models—using real-world driving data drawn from two scenarios defined in United Nations Regulation No. 157. Our evaluation on approximately 18 million instances of these two scenarios demonstrates that Gaussian Mixture Copula Models consistently surpass Gaussian Copula Models and perform better than, or at least comparably to, Gaussian Mixture Models, as measured by both log-likelihood and Sinkhorn distance. These results are promising for the adoption of Gaussian Mixture Copula Models as a statistical foundation for future scenario-based validation frameworks.
@@ 1 INTRODUCTION
  • Gaussian Mixture Models (GMMs) approximate the joint density with a weighted sum of elliptical Gaussians [14] and have been successfully applied to driving scenarios, such as the highway cut-in risk analysis by Akagi et al. [15]. While GMMs can approximate almost any continuous distribution given a sufficiently large number of components [16], practical applications typically rely on finite mixtures with a constrained number of components. In this finite regime, their reliance on multivariate Gaussian components can be limiting in practice when capturing intricate dependencies between parameters: the Gaussian assumptions may oversimplify complex relationships and higher-order interactions, thereby limiting their ability to model non-linear dependencies effectively.
@@ A GMM describes the joint density of random variables as a weighted sum of $K$ normal distributions,
where $\alpha^{k}\geq 0$ denotes the weight of the $k$-th component, such that $\sum_{k=1}^{K}\alpha^{k}=1$. $\phi(x_{1},x_{2},\dots x_{d};\bm{\mu}^{k},\Sigma^{k})$ represents a multivariate normal distribution with mean vector $\bm{\mu}^{k}$ and covariance matrix $\Sigma^{k}$. The parameters $\Theta=\{\alpha^{k},\bm{\mu}^{k},\Sigma^{k}\}_{k=1}^{K}$ of a GMM can be efficiently learned using the expectation-maximization (EM) algorithm [24]. The two-step approach of the EM algorithm guarantees convergence to a local maximum of the likelihood function. As stated before, a limitation of GMMs is the assumption that each of the $K$ components follows a multivariate normal distribution. For finite mixtures, this assumption can become restrictive in practice when the data exhibit non-Gaussian characteristics, such as skewness, heavy-tailed distributions, or intricate dependencies between variables, as is the case for our scenario parameters.
@@ 3 ESTIMATION OF GMCM PARAMETERS
Estimating the dependencies among the random variables, represented by the GMC parameters $\Theta=\left\{\alpha^{k},\mkern 1.0mu\bm{\mu}^{k},\mkern 1.0mu\Sigma^{k}\right\}_{k=1}^{K}$ is non-trivial. It is neither possible to directly determine these from an empirical correlation matrix, as is the case for Gaussian copulas, nor is the EM algorithm for GMM parameter estimation directly applicable. The (log-)likelihood function (7) of GMCMs does not admit a closed-form analytical expression, in particular because of the function $\Psi^{-1}$ appearing in it. Further, GMCMs suffer from an inherent issue of parameter non-identifiability. These challenges and structural properties have been previously examined by Tewari et al. [19, 23].
@@ 4 EXPERIMENT
In this section, we present our experimental study of GMCMs for joint probability estimation of scenario parameters. Our experiments, conducted on two real-world datasets from certification-relevant scenarios, show that GMCMs consistently outperform GCMs. With comparable computational resources, GMCMs achieve higher accuracy than GMMs for one scenario and comparable performance for the other. These results confirm previously reported advantages of GMCMs [19, 23] and extend them to the domain of scenario parameter modeling.
@@ 4.1 REAL-WORLD SCENARIO DATASETS
The scenario Deceleration is provided as one of three traffic critical disturbance scenarios to define conditions under which ALKS shall avoid a collision. The scenario conditions (i.e., parameter values) are determined by simulating the scenarios with an attentive human-driver performance model.
@@ 4.2 RESULTS
To illustrate the achieved expressive power of the GMCM, we visualize the bivariate marginal distribution of LeadVehicle Init HeadwayTime (s) and Ego InitSpeed Ve0 (kph) from the Follow Lead Vehicle Emergency Brake dataset via density heatmaps. Fig. 4 compares the original data distribution—which appears multimodal with non-Gaussian modes—with the estimated densities from the GMCM, the GMM, and the GCM.
In the high-density region on the left, all three models deviate from the empirical distribution, but both GMCM and GMM still capture the main structure with a descending trend of LeadVehicle Init HeadwayTime (s) towards higher Ego InitSpeed Ve0 (kph), whereas the GCM tends to smooth out this pattern. In the central high-density region and the right part of the plot, the GMCM more faithfully reproduces the location and shape of the empirical distribution. In contrast, the GMM smooths out the central high-density region, and the GCM assigns too much density to the upper-right part of the plot. Overall, the visual comparison of the bivariate density heatmaps suggests that the GMCM provides the closest match to the real data distribution.